# Loss Landscape Degeneracy Drives Stagewise Development in Transformers

### v2

@@ 5 Degeneracy-based stage division
$F_{n}(W^{*})=n\ell_{n}({w^{*}})+\lambda({w^{*}})\log n-(m({w^{*}})-1)\log\log n+O_{p}(1).$ (4)
Here, $\ell_{n}({w^{*}})$ is the empirical loss, $\lambda({w^{*}})$ is the LLC, $m({w^{*}})$ is the local multiplicity, and the lower-order terms include a constant contribution from the prior mass of $W^{*}$.
The first two terms in equation 4 create a tradeoff between accuracy ($\ell_{n}$) and degeneracy ($\lambda$). Moreover, as $n$ increases, the linear term becomes increasingly important relative to the logarithmic term, changing the nature of the tradeoff. At certain critical $n$ the neighborhood with the lowest local free energy may rapidly change to a neighborhood with decreased loss and increased LLC, as illustrated in Figure 3.
@@
This perspective suggests that changes in the loss landscape degeneracy, as measured by the LLC, reflect qualitative changes in the model. In larger models, we expect that these qualitative changes may be more gradual, while still being delineated by brief moments in which the posterior is stably concentrated around a given local minimum. This motivates our approach of identifying plateaus in the estimated LLC curve—brief pauses before and after a given increase or decrease in degeneracy—as stage boundaries which divide training into developmental stages.
@@ Results
In our experiments, we identify plateaus in the estimated LLC curve by first lightly smoothing the LLC curve with a Gaussian process to facilitate stable numerical differentiation with respect to log training time. We identify plateaus as approximate zeros of this derivative, namely local minima of the absolute derivative that fall below a small threshold (see Section B.1). Figures 1, B.2 and B.3 show the results. Sections B.4 and B.4 shows that similar stage divisions arise for independent training runs.
@@
6 Results for language modeling
Plateaus in LLC estimates (Figures 1(a) and 1(a)) reveal five developmental stages for our language model. In order to validate that this stage division is meaningful, we search for concomitant changes in the model’s input/output behavior and its internal computational structure. In this section, we report a range of setting-specific metrics that reveal the following significant, interpretable changes coinciding with each stage: in LM1 the model learns to predict according to bigram statistics; in LM2 the model learns to predict frequent $n$-grams and use the positional embedding; in LM3 and LM4 the model respectively forms “previous-token heads” and “induction heads” as part of the same induction circuit studied by Olsson et al. (2022). Note that we did not discover significant changes in LM5, and we do not claim that these are the only interesting developmental changes occurring throughout training. There may be other interesting developmental changes that are not captured by our metrics, or are not significant enough to not show up in the LLC curve.
6.1 Stage LM1 (0–900 steps)
Learning bigram statistics
Figure 4(a) shows that the bigram score—the average cross entropy between model logits and empirical bigram frequencies (see Section C.1.1)—is minimized around the LM1LM2 boundary, with a value only 0.3 nats above the irreducible entropy of the empirical bigram distribution. This suggests that during LM1 the model learns to predict using bigram statistics (the optimal next-token prediction given only the current token).
6.2 Stage LM2 (900–6.5k steps)
Using positional information
During LM2 the positional embedding becomes structurally important. Figure 4(b) shows that here the test loss for the model with the positional embedding zero-ablated diverges from the test loss of the unablated model (see Section C.2.1). There is also an uptick in previous-token attention among some first-layer attention heads shown in green in Figure 4(d).
Learning common $n$-grams
We define an $n$-gram score as the ratio of final-position token loss on (1) a baseline set of samples from a validation set truncated to $n$ tokens, and (2) a fixed set of common $n$-grams (see Section C.1.2). Figure 4(c) shows a large improvement in $n$-gram score for $n=3,4$ during LM2. This suggests that during LM2 the model memorizes and learns to predict common $n$-grams for $n>2$ (note this requires using the positional encoding and may also involve previous-token heads).
@@ 6.1 Stage LM1 (0–900 steps)
In this stage, the heads that eventually become previous-token and induction heads in future stages begin to compose (that is, read from and write to a shared residual stream subspace; see Figures C.4 and C.2.2). This suggests that the foundations for the induction circuit are laid in advance of any measurable change in model outputs or attention patterns.
6.3 Stages LM3, LM4 (6.5k–8.5k, 8.5k–17k steps)
@@
Figure 4(d) shows the previous-token matching score (Section C.2.3) rises over LM3 and LM4 for the two first-layer heads that eventually participate in the induction circuit (as distinguished by their composition scores, Section C.2.2). Figure 4(e) shows that during LM4 there is an increase in the prefix-matching score (Section C.2.4) for the two second-layer induction heads that complete the induction circuit. Figure 4(f) shows a corresponding drop in the ICL score (Section C.1.3) as the model begins to perform in-context learning.
@@ 7 Results for in-context linear regression
Plateaus in the LLC estimates (Figures 1(b) and 1(b)) reveal five developmental stages for our in-context linear regression model. We validate that this stage division is meaningful by identifying significant, concomitant changes in the model’s structure and behavior: in LR1 the model learns to predict without looking at the context; in LR2 the model acquires a robust in-context learning ability; and in LR3 and LR4 the model becomes more fragile to out-of-distribution inputs. We did not discover significant changes in LR5, nor do we claim this is an exhaustive list of developments.
@@ 7.1 Stage LR1 (0–1k steps)
Figure 5(a) shows that the mean square prediction for all tokens $\mathbb{E}[\|\hat{y}_{k}\|^{2}]$ decreases during LR1, reaching a minimum of $0.1$ (smaller than the target noise $\sigma^{2}=0.125$) slightly after the end of LR1. Similar to how the language model learned bigram statistics in LM1, this suggests the model first learns the optimal context-independent prediction $\hat{y}_{k}=\bar{\mathbf{t}}^{\top}x_{k}$ where $\bar{\mathbf{t}}$ is the mean of the task distribution (zero in this case).
@@ 7.2 Stage LR2 (1k–40k steps)
Figure 5(b) shows that during LR2 there is a drop in ICL score (Section D.1.2), indicating that the model acquires in-context learning.
@@
Section D.2 documents additional changes. Near the end of LR2, token and positional embeddings begin to “collapse,” effectively losing singular values and aligning with the same activation subspace (Sections D.2.1 and D.2.2). At the same time, several attention heads form concentrated, input-independent attention patterns (Section D.2.3).
7.3 Stages LR3, LR4 (40k–126k, 126k–320k steps)
@@
While performance continues to improve on typical sequences, Figure 5(b) shows that during LR3 and LR4, the model’s in-context learning ability deteriorates for outlier sequences with higher-than-average $|x_{k}|$.
@@
Figure 5(c) shows the individual weights in the final layer normalization module. A large fraction of these weights go to zero in LR3 and LR4. This occurs in tandem with a similar collapse in the weights of the unembedding transforms (Section D.2.4). This results in the model learning to read its prediction $\hat{y}_{k}$ from a handful of privileged dimensions of the residual stream. Since this means that the network outputs become insensitive to changes in many of the parameters, we conjecture that this explains part of the striking decrease in estimated LLC over these stages (Section D.2.4).
This collapse is most pronounced and affects the largest proportion of weights in the unembedding module, but in LR4 it spreads to earlier layer normalization modules, particularly the layer normalization module before the first attention block (Section D.2.5).
@@ 8 Discussion
While we have described the LLC as a measure of loss landscape degeneracy, it can also be understood as a measure of model complexity (cf. Section A.1). It is natural for changes in a model’s internal structure to show up as a change in complexity. For example, Chen et al. (2024) showed that the emergence of syntactic attention structure coincides with a spike in two model complexity measures, namely the model’s Fisher information and the intrinsic dimension (Facco et al., 2017) of the model’s embeddings.
@@
This phenomenon is currently not accounted for by theories of neural network development. In the theory of saddle-to-saddle dynamics, deep linear networks learn progressively more complex approximations of the data (Saxe et al., 2019). Likewise, the example transitions in the singular learning process outlined in Sections 5 and 3 describe LLC increases. Though we note that decreasing the LLC while holding the loss constant would be another way to decrease the free energy according to equation 4, providing a full theoretical account of these stages is an open problem.
Contribution statement
@@ Acknowledgments
Appendix A reviews the learning coefficient, providing some simple toy examples contrasting the learning coefficient with Hessian-based measures. This section also discusses SGLD-based LLC estimation including experiment hyperparameters (Section A.3), and offers a detailed example of the calibrations involved in applying LLC estimation to regression transformers to serve as a reference (Section A.4). Appendix B provides further detail on our procedure for LLC-based stage identification, including stages identified in additional training runs and a brief comparison with Hessian statistics. Appendices C and D examine the developmental stages of language models and in-context linear regression in more detail and explain the various metrics we use to track behavioral and structural development. Appendix E describes some additional experiments on a one-layer language model. Appendix F covers transformer training experimental details, such as model architectures, training procedures, and hyperparameters.
@@ Appendix A The local learning coefficient (LLC)
A.1 Interpretations and examples of the LLC
In Section 4, we introduced the LLC as a quantification of geometric degeneracy. In this section, we discuss an additional perspectives on the LLC as a count of the “effective” dimensionality of a parameter, and we give additional examples of the LLC. We refer the reader to Watanabe (2009) and Lau et al. (2025) for more discussion.
@@ A.1 Interpretations and examples of the LLC
Figure 2 offers several additional examples, from left to right:
@@
A.2 Estimating LLCs with SGLD
@@
In each step $\tau$ we sample a mini-batch of size $m$ and the associated empirical loss, denoted $\ell_{m}^{(\tau)}$, is used to compute the gradient in the SGLD update. We note that LLC estimator defined in 3 uses the expectation $\mathbb{E}^{\beta}[\ell_{n}(w)]$ which in the current notation means we should take $\mathcal{O}(w)$ to be $\ell_{n}(w)$. For computational efficiency we follow Lau et al. (2025) in recycling the mini-batch losses $\ell_{m}(w^{(c)}_{\tau})$ computed during the SGLD process. That is, we take $\mathcal{O}=\ell_{m}^{(\tau)}$ rather than $\mathcal{O}=\ell_{n}$.
A.3 LLC estimation experiment details
A.3.1 LLC estimation details for language models
For language models, we use SGLD to sample 20 independent chains with 200 steps per chain and 1 sample per step, at a temperature $\beta=1/\log(m)$, where $m=100$ is the size of the batch (the maximum size that would fit in memory). For the one-layer model, we used $\epsilon=0.003,\gamma=300$, and for the two-layer model we used $\epsilon=0.001,\gamma=100$. Estimating the LLC across all checkpoints took around 200 GPU hours for the two-layer model on a single A100 and around 125 GPU hours for the one-layer model. For additional runs of the two-layer model, we ran fewer chains, bringing the time down to about 2 TPU hours per training run.
@@ A.3.1 LLC estimation details for language models
C Sampler # of chains $20$
$T_{\mathrm{SGLD}}$ Sampler # of SGLD steps / chain $200$
$\epsilon$ SGLD Step size $0.003$ $0.001$
$\tilde{\gamma}=\epsilon\gamma/2$ SGLD Localization strength $300$ $200$
$\tilde{\beta}=\epsilon\beta/2n$ SGLD Inverse temperature 0.0000217147
$m$ SGLD (Default: $\beta^{*}=\frac{1}{\log n}$) 100
The size of each SGLD batch
@@ C Sampler # of chains $20$
A.3.2 LLC estimation details for in-context linear regression
For in-context linear regression models, we generate a fixed dataset of $2^{20}$ samples. Using SGLD, we sample $10$ independent chains with 5,000 steps per chain, of which the first 1,000 are discarded as a burn-in, after which we draw observations once per step, at a temperature $\epsilon\beta/2n=0.01$, $\epsilon=0.0003$, and $\epsilon\gamma/2=0.01$, over batches of size $m=1024$. LLC estimation takes up to 72 CPU-hours per training run.

### v3

@@ 5 Degeneracy-based stage division
$F_{n}(W^{*})=n\ell_{n}({w^{*}})+\lambda({w^{*}})\log n+O_{p}(\log\log n)$ (4)
Here, $\ell_{n}({w^{*}})$ is the empirical loss, $\lambda({w^{*}})$ is the LLC, and the lower-order terms include a constant contribution from the prior mass of $W^{*}$.
The first two terms in equation ˜4 create a tradeoff between accuracy ($\ell_{n}$) and degeneracy ($\lambda$). Moreover, as $n$ increases, the linear term becomes increasingly important relative to the logarithmic term, changing the nature of the tradeoff. At certain critical $n$ the neighborhood with the lowest local free energy may rapidly change to a neighborhood with decreased loss and increased LLC, as illustrated in Figure˜3.
@@
This perspective suggests that changes in the loss landscape degeneracy, as measured by the LLC, reflect qualitative changes in the model. In larger models, we expect that these qualitative changes may be more gradual, while still being delineated by brief moments in which the posterior is stably concentrated around a given local minimum. This motivates our approach of identifying plateaus in the estimated LLC curve—brief pauses before and after a given increase or decrease in degeneracy—as stage boundaries which divide training into approximate developmental stages. The resolution of these stage boundaries depends on the density of checkpoints used for LLC estimation and the precision of those estimates.
@@ Results
In our experiments, we identify plateaus in the estimated LLC curve by first lightly smoothing the LLC curve with a Gaussian process to facilitate stable numerical differentiation with respect to log training time. We identify plateaus as approximate zeros of this derivative, namely local minima of the absolute derivative that fall below a small threshold (see Section˜B.1). Figures˜1, B.2 and B.3 show the results. Sections˜B.4 and B.4 shows that similar stage divisions arise for independent training runs.
6 Results for language modeling
Plateaus in LLC estimates (Figures˜1(a) and 1(a)) reveal five developmental stages for our language model. In order to validate that this stage division is meaningful, we search for concomitant changes in the model’s input/output behavior and its internal computational structure. In this section, we report a range of setting-specific metrics that reveal the following significant, interpretable changes coinciding with each stage: in LM1 the model learns to predict according to bigram statistics; in LM2 the model learns to predict frequent $n$-grams and use the positional embedding; in LM3 and LM4 the model respectively forms “previous-token heads” and “induction heads” as part of the same induction circuit studied by Olsson et al. (2022). Note that we did not discover significant changes in LM5, and we do not claim that these are the only interesting developmental changes occurring throughout training. There may be other interesting developmental changes that are not captured by our metrics, or are not significant enough to not show up in the LLC curve.
6.1 Stage LM1 (0–900 steps)
Learning bigram statistics
Figure˜4(a) shows that the bigram score—the average cross entropy between model logits and empirical bigram frequencies (see Section˜C.1.1)—is minimized around the LM1LM2 boundary, with a value only 0.3 nats above the irreducible entropy of the empirical bigram distribution. This suggests that during LM1 the model learns to predict using bigram statistics (the optimal next-token prediction given only the current token).
6.2 Stage LM2 (900–6.5k steps)
Using positional information
During LM2 the positional embedding becomes structurally important. Figure˜4(b) shows that here the test loss for the model with the positional embedding zero-ablated diverges from the test loss of the unablated model (see Section˜C.2.1). Specifically, we mean setting learned positional embeddings to zero during evaluation. Conditional on our architecture this establishes whether the model effectively uses positional information. A similar method could be used in a model without learned positional embeddings. There is also an uptick in previous-token attention among some first-layer attention heads shown in green in Figure˜4(d).
Learning common $n$-grams
We define an $n$-gram score as the ratio of final-position token loss on (1) a baseline set of samples from a validation set truncated to $n$ tokens, and (2) a fixed set of common $n$-grams (see Section˜C.1.2). To compute the “common n-grams” score after extracting the top 1000 n-grams, we compute the loss on contexts like
            [<bos_token>, <token_1>, <token_2>, ..., <token_n>]
using the loss on <token_n> and normalize against the average loss on the $n$-th token of similar-length contexts drawn from the pretraining distribution, then divide the $n$-th token loss of truncated pretraining contexts by the $n$-gram loss to get the $n$-gram score.
Figure˜4(c) shows a large improvement in $n$-gram score for $n=3,4$ during LM2. This suggests that during LM2 the model memorizes and learns to predict common $n$-grams for $n>2$ (note this requires using the positional encoding and may also involve previous-token heads).
@@ 6.1 Stage LM1 (0–900 steps)
In this stage, the heads that eventually become previous-token and induction heads in future stages begin to compose (that is, read from and write to a shared residual stream subspace; see Figures˜C.4 and C.2.2). This suggests that the foundations for the induction circuit are laid in advance of any measurable change in model outputs or attention patterns.
6.3 Stages LM3 & LM4 (6.5k–8.5k & 8.5k–17k steps)
@@
Figure˜4(d) shows the previous-token matching score (Section˜C.2.3) rises over LM3 and LM4 for the two first-layer heads that eventually participate in the induction circuit (as distinguished by their composition scores, Section˜C.2.2). Figure˜4(e) shows that during LM4 there is an increase in the prefix-matching score (Section˜C.2.4) for the two second-layer induction heads that complete the induction circuit. Figure˜4(f) shows a corresponding drop in the ICL score (Section˜C.1.3) as the model begins to perform in-context learning.
@@ 7 Results for in-context linear regression
Plateaus in the LLC estimates (Figures˜1(b) and 1(b)) reveal five developmental stages for our in-context linear regression model. We validate that this stage division is meaningful by identifying significant, concomitant changes in the model’s structure and behavior: in LR1 the model learns to predict without looking at the context; in LR2 the model acquires a robust in-context learning ability; and in LR3 and LR4 the model becomes more fragile to out-of-distribution inputs. We did not discover significant changes in LR5, nor do we claim this is an exhaustive list of developments.
@@ 7.1 Stage LR1 (0–1k steps)
Figure˜5(a) shows that the mean square prediction for all tokens $\mathbb{E}[\|\hat{y}_{k}\|^{2}]$ decreases during LR1, reaching a minimum of 0.10.10.1 (smaller than the target noise $\sigma^{2}=0.125$) slightly after the end of LR1. Similar to how the language model learned bigram statistics in LM1, this suggests the model first learns the optimal context-independent prediction $\hat{y}_{k}=\bar{}\mathbf{t}^{\top}x_{k}$ where ¯$\bar{}\mathbf{t}$ is the mean of the task distribution (zero in this case).
@@ 7.2 Stage LR2 (1k–40k steps)
Figure˜5(b) shows that during LR2 there is a drop in ICL score (Section˜D.1.2), indicating that the model acquires in-context learning.
@@
Section˜D.2 documents additional changes. Near the end of LR2, token and positional embeddings begin to “collapse,” effectively losing singular values and aligning with the same activation subspace (Sections˜D.2.1 and D.2.2). At the same time, several attention heads form concentrated, input-independent attention patterns (Section˜D.2.3).
7.3 Stages LR3 & LR4 (40k–126k & 126k–320k steps)
@@
While performance continues to improve on typical sequences, Figure˜5(b) shows that during LR3 and LR4, the model’s in-context learning ability deteriorates for outlier sequences with higher-than-average $|x_{k}|$.
@@
Figure˜5(c) shows the individual weights in the final layer normalization module. A large fraction of these weights go to zero in LR3 and LR4. This occurs in tandem with a similar collapse in the weights of the unembedding transforms (Section˜D.2.4). This results in the model learning to read its prediction $\hat{y}_{k}$ from a handful of privileged dimensions of the residual stream. Since this means that the network outputs become insensitive to changes in many of the parameters, we conjecture that this explains part of the striking decrease in estimated LLC over these stages (Section˜D.2.4).
This collapse is most pronounced and affects the largest proportion of weights in the unembedding module, but in LR4 it spreads to earlier layer normalization modules, particularly the layer normalization module before the first attention block (Section˜D.2.5).
@@ 8 Discussion
While we have described the LLC as a measure of loss landscape degeneracy, it can also be understood as a measure of model complexity (cf. Section˜A.2). It is natural for changes in a model’s internal structure to show up as a change in complexity. For example, Chen et al. (2024) showed that the emergence of syntactic attention structure coincides with a spike in two model complexity measures, namely the model’s Fisher information and the intrinsic dimension (Facco et al., 2017) of the model’s embeddings.
@@
This phenomenon is currently not accounted for by theories of neural network development. In the theory of saddle-to-saddle dynamics, deep linear networks learn progressively more complex approximations of the data (Saxe et al., 2019). Likewise, the example transitions in the singular learning process outlined in Sections˜5 and 3 describe LLC increases. Though we note that decreasing the LLC while holding the loss constant would be another way to decrease the free energy according to equation ˜4, providing a full theoretical account of these stages is an open problem.
Author contributions
@@ Acknowledgments
Appendix˜A reviews the learning coefficient, providing some simple toy examples contrasting the learning coefficient with Hessian-based measures. This section also discusses SGLD-based LLC estimation including experiment hyperparameters (Section˜A.4), and offers a detailed example of the calibrations involved in applying LLC estimation to regression transformers to serve as a reference (Section˜A.5). Appendix˜B provides further detail on our procedure for LLC-based stage identification, including stages identified in additional training runs and a brief comparison with Hessian statistics. Appendices˜C and D examine the developmental stages of language models and in-context linear regression in more detail and explain the various metrics we use to track behavioral and structural development. Appendix˜E describes some additional experiments on a one-layer language model. Appendix˜F covers transformer training experimental details, such as model architectures, training procedures, and hyperparameters.
@@ Appendix A The local learning coefficient (LLC)
A.1 Formal Definition of the LLC
In the setting of Section˜4, let $B$ be a closed ball around ${w^{*}}$ such that ${w^{*}}$ is a global minimum on $B$, by which we mean a point with (equal) lowest loss. If there are multiple such global minima, the volume asymptotics are determined by the geometry of one that is most degenerate in the precise sense of SLT, formalised in Lau et al. (2025), roughly corresponding to having the lowest LLC. We call this minimum the maximally degenerate global minimum on $B$. Consider the volume of the set of nearby low-loss parameters,
[…]
As $\epsilon\to 0$, $V(\epsilon)$ is asymptotically equivalent to
$c\epsilon^{\lambda({w^{*}})}\log(1/\epsilon)^{m({w^{*}})-1},$
where $\lambda({w^{*}})$ is the LLC, $m({w^{*}})$ is another geometric quantity called the local multiplicity, and $c>0$ is a constant.
A.2 Interpretations and examples of the LLC
In Section˜4, we introduced the LLC as a quantification of geometric degeneracy. In this section, we discuss an additional perspectives on the LLC as a count of the “effective” dimensionality of a parameter, and we give additional examples of the LLC. We refer the reader to Watanabe (2009) and Lau et al. (2025) for more discussion.
@@ A.2 Interpretations and examples of the LLC
Figure˜2 offers several additional examples, from left to right:
@@
A.3 Estimating LLCs with SGLD
@@
In each step $\tau$ we sample a mini-batch of size $m$ and the associated empirical loss, denoted $\ell_{m}^{(\tau)}$, is used to compute the gradient in the SGLD update. We note that LLC estimator defined in ˜3 uses the expectation $\mathbb{E}^{\beta}[\ell_{n}(w)]$ which in the current notation means we should take $\mathcal{O}(w)$ to be $\ell_{n}(w)$. For computational efficiency we follow Lau et al. (2025) in recycling the mini-batch losses $\ell_{m}(w^{(c)}_{\tau})$ computed during the SGLD process. That is, we take $\mathcal{O}=\ell_{m}^{(\tau)}$ rather than $\mathcal{O}=\ell_{n}$.
Time and Space Complexity.
The computational cost per LLC estimate is proportional to a standard training step, denoted $S$. We expect to require a constant number $CT_{\text{SGLD}}$ of samples (on the order of $10^{2}$$10^{4}$) to yield robust estimates, independent of model size. Using logarithmically spaced checkpoints, the total computational complexity for generating an LLC curve over the entire training process scales as $O(SCT_{\text{SGLD}}\log N)$, where $N$ is the total number of training steps. The space complexity incurs a modest linear overhead compared to standard SGD, requiring storage for one additional copy of the weights to enable localization.
A.4 LLC estimation experiment details
A.4.1 LLC estimation details for language models
For language models, we use SGLD to sample 20 independent chains with 200 steps per chain and 1 sample per step. For the one-layer model, we used $\epsilon=0.003,\gamma=300$, and for the two-layer model we used $\epsilon=0.001,\gamma=100$. Estimating the LLC across all checkpoints took around 200 GPU hours for the two-layer model on a single A100 and around 125 GPU hours for the one-layer model. For additional runs of the two-layer model, we ran fewer chains, bringing the time down to about 2 TPU hours per training run.
@@ A.4.1 LLC estimation details for language models
C Sampler # of chains 202020
$T_{\mathrm{SGLD}}$ Sampler # of SGLD steps / chain 200200200
$\epsilon$ SGLD Step size 0.0030.0030.003 0.0010.0010.001
$\gamma$ SGLD Localization strength 300 100
$n\beta$ SGLD Inverse temperature 21.7
$m$ SGLD The size of each SGLD batch 100
@@ C Sampler # of chains 202020
A.4.2 LLC estimation details for in-context linear regression
For in-context linear regression models, we generate a fixed dataset of $2^{20}$ samples. Using SGLD, we sample 101010 independent chains with 5,000 steps per chain, of which the first 1,000 are discarded as a burn-in, after which we draw observations once per step, at a temperature $n\beta=66.7$, $\epsilon=0.0003$, and $\gamma=13.3$, over batches of size $m=1024$. LLC estimation takes up to 72 CPU-hours per training run.